OpenStudy (anonymous):

find the product (8x^2-5y)(8x^2+5y) can someone help

OpenStudy (callisto):

Use the identity (a-b)(a+b) = a^2 - b^2 Now, put a= 8x^2 , b=5y into the identity (8x^2-5y)(8x^2+5y) = (8x^2)^2 - (5y)^2 = ?